# EvoGPT: Leveraging LLM-Driven Seed Diversity to Improve Search-Based Test Suite Generation

### v2

@@ III-B2 Selection
EvoGPT employs ranked selection to choose parent test suites for reproduction. The population is sorted by our fitness function, and selection probabilities are assigned according to rank rather than raw fitness values. Ranked selection is a common choice in EAs [whitley1989genitor, goldberg1991comparative, baker2014adaptive], as it helps control selection pressure and reduce the risk of premature convergence without requiring problem-specific tuning.
@@ III-B4 Mutation
An additional operator used in EAs is Mutation. Mutation introduces random variations into individuals, ensuring continued diversity and helping escape local optima. We implement mutation operators based on those used by EvoSuite [fraser2012evosuite], applying structural transformations to test methods. Our mutation strategy employs three operators, each selected uniformly at random:
@@ III-B5 Plateau Escape via LLM Injection
EAs can become trapped in local optima or plateau, where standard genetic operators fail to significantly improve fitness despite continued iterations. Inspired by CodaMosa [lemieux2023codamosa], EvoGPT incorporates an LLM Injection mechanism to escape such plateaus. The five LLM agents mentioned in previous section are queried in parallel. Each agent receives as input the current test suite context (imports, fields, and setup methods), the source code under test, and detailed coverage feedback including the specific uncovered branches and mutants identified. Rather than generating complete test suites, each LLM agent only adds new test methods within an existing test suite (i.e., only @Test methods). These new tests are validated and injected into the current best individual in the population. This design allows EvoGPT to enrich promising test suites with targeted behaviors without disrupting their existing structure, enabling the EA to escape local optima when stagnating, and resume effective exploration.
@@ IV Experimental Results
We conducted an experimental evaluation of EvoGPT using Defects4J [just2014defects4j], a benchmark of real-world Java projects with reproducible bugs and accompanying test suites. Defects4J is commonly used in the automated testing literature due to its diversity, complexity, and grounding in real software engineering practice. Table II presents the Defects4J projects used in our experiments. Columns “Identifier”, “Project name”, “Version”, and “Focals” represent the identifier of the repository under test, its name, the version of the code being tested, and the number of focal methods extracted from each project, respectively. Overall we experimented with all 17 repositories in the Defects4J benchmark and extracted at least 137 focal methods for each repository.
@@
We compared EvoGPT to two baselines: EvoSuite [fraser2012evosuite], a widely used EA-based test generation system, and TestART [2024arXiv240803095G], a recent state-of-the-art system for LLM-based unit test generation. We chose TestART since it has been shown to outperform previous LLM-based approaches such as ChatUniTest [chen2024chatunitest]. We did not include hybrid LLM-SBST approaches as baselines in our experiments as CodaMosa [lemieux2023codamosa] and pytLMtester [yang2025llm] are specifically designed for Python, a dynamically typed language, and SearchSYS [dakhama2025enhancing] targets system-level simulators rather than unit-level Java programs.
@@ IV-A Comparison Against Baselines
Next, we conducted Wilcoxon signed-rank tests [wilcoxon1992individual] to validate that the results above are statistically significant. We compared EvoGPT against TestART and EvoSuite across all of our dataset projects ($n=17$ projects) and evaluation metrics. We chose the Wilcoxon test as it is a non-parametric alternative appropriate for paired samples without assuming normal distribution, following recommendations for empirical software engineering studies [arcuri2011practical].
The tests were performed over per-project results for LCCT, BCCT and MSCT across all 17 benchmark projects. We report also Cliff’s delta ($\delta$), which is the ratio of cases where one algorithm was better than the other minus the ratio of the cases where the other algorithm was better [cliff1993dominance]. This is a known way to measure effect size, where $|\delta|\geq 0.474$ indicates a large effect.

### v3

@@ III-B2 Selection
EvoGPT employs ranked selection to choose parent test suites for reproduction. The population is sorted by our fitness function, and selection probabilities are assigned according to rank rather than raw fitness values. Ranked selection is a common choice in EAs [31, 17, 6], as it helps control selection pressure and reduce the risk of premature convergence without requiring problem-specific tuning.
@@ III-B4 Mutation
An additional operator used in EAs is Mutation. Mutation introduces random variations into individuals, ensuring continued diversity and helping escape local optima. We implement mutation operators based on those used by EvoSuite [15], applying structural transformations to test methods. Our mutation strategy employs three operators, each selected uniformly at random:
@@ III-B5 Plateau Escape via LLM Injection
EAs can become trapped in local optima or plateau, where standard genetic operators fail to significantly improve fitness despite continued iterations. Inspired by CodaMosa [23], EvoGPT incorporates an LLM Injection mechanism to escape such plateaus. The five LLM agents mentioned in previous section are queried in parallel. Each agent receives as input the current test suite context (imports, fields, and setup methods), the source code under test, and detailed coverage feedback including the specific uncovered branches and mutants identified. Rather than generating complete test suites, each LLM agent only adds new test methods within an existing test suite (i.e., only @Test methods). These new tests are validated and injected into the current best individual in the population. This design allows EvoGPT to enrich promising test suites with targeted behaviors without disrupting their existing structure, enabling the EA to escape local optima when stagnating, and resume effective exploration.
@@ IV Experimental Results
We conducted an experimental evaluation of EvoGPT using Defects4J [22], a benchmark of real-world Java projects with reproducible bugs and accompanying test suites. Defects4J is commonly used in the automated testing literature due to its diversity, complexity, and grounding in real software engineering practice. Table II presents the Defects4J projects used in our experiments. Columns “Identifier”, “Project name”, “Version”, and “Focals” represent the identifier of the repository under test, its name, the version of the code being tested, and the number of focal methods extracted from each project, respectively. Overall we experimented with all 17 repositories in the Defects4J benchmark and extracted at least 137 focal methods for each repository.
@@
We compared EvoGPT to two baselines: EvoSuite [15], a widely used EA-based test generation system, and TestART [18], a recent state-of-the-art system for LLM-based unit test generation. We chose TestART since it has been shown to outperform previous LLM-based approaches such as ChatUniTest [12]. We did not include hybrid LLM-SBST approaches as baselines in our experiments as CodaMosa [23] and pytLMtester [35] are specifically designed for Python, a dynamically typed language, and SearchSYS [14] targets system-level simulators rather than unit-level Java programs.
@@ IV-A Comparison Against Baselines
Next, we conducted Wilcoxon signed-rank tests [32] to validate that the results above are statistically significant. We compared EvoGPT against TestART and EvoSuite across all of our dataset projects ($n=17$ projects) and evaluation metrics. We chose the Wilcoxon test as it is a non-parametric alternative appropriate for paired samples without assuming normal distribution, following recommendations for empirical software engineering studies [5].
The tests were performed over per-project results for LCCT, BCCT and MSCT across all 17 benchmark projects. We report also Cliff’s delta ($\delta$), which is the ratio of cases where one algorithm was better than the other minus the ratio of the cases where the other algorithm was better [13]. This is a known way to measure effect size, where $|\delta|\geq 0.474$ indicates a large effect.